OpenStudy (anonymous):

How to use the substitution method in this integral? ∫sin (x^3 - 2X +1)(3X^2 -2)dX

OpenStudy (agreene):

Let u= x^3-2x+1 du = 3x^2-2 Which gives us: \[\int \sin u du\]